# Who’s Asking? Investigating Bias Through the Lens of Disability-Framed Queries in LLMs

### v1

@@ 2.1 Bias in LLMs
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of NLP tasks; however, they remain susceptible to representational bias and stereotyping. Prior research has shown that LLMs often absorb and reproduce harmful societal biases - especially those tied to gender, race, religion, and health - across both contextual embeddings and generative outputs [22, 9, 18, 1, 28]. These models frequently reflect the imbalances present in their training data, reproducing gendered, socioeconomic, and occupational assumptions rather than equitable or contextually informed representations [13, 24, 17, 27]. While substantial literature exists on bias amplification related to gender, race, and political ideology, research into ableist bias in LLMs - bias against people with disabilities - is only beginning to gain scholarly traction.
@@ 2.2 Disability bias
Emerging studies [8, 15, 35] reveal that LLMs often encode and reflect deeply entrenched stereotypes about people with disabilities. For instance, models tend to associate disability-related terms with negative sentiment [14], echoing societal stigma found in training data. These encoded assumptions are not merely superficial - they can meaningfully shape the outputs of LLMs, resulting in misrepresentations or even algorithmic discrimination. This is particularly concerning in high-stakes contexts like hiring [15], where disclosure of a disability can lead to skewed responses. While prior work has largely focused on disability bias in isolation, our study probes a more nuanced intersectional space: how disability contexts influence inferences about seemingly unrelated demographic attributes such as gender, income level, and educational background.
@@ 2.4 Intersectionality of biases
There remains limited but growing attention on how intersecting identities influence LLM behavior. While several recent works [6, 14, 19, 20] have evaluated how multiple marginalized identities interact within prompts, few studies have examined how disability cues lead to inferred biases about unrelated attributes. Notably, [19] explored ableist bias in tandem with other identity categories, using sentiment and social perception metrics. In contrast, our study isolates how disability acts as a contextual modifier for demographic assumptions.
To support such analysis, various benchmark datasets have emerged. For instance, [6] introduced a probing framework for identity-based behavior using persona-based prompts, while [20] presented a stereotype prevalence dataset for intersectional identities across LLMs. These newer datasets emphasize richer, sentence-level stimuli that better reflect real-world complexity [5].
@@ 3.1 Dataset Selection
For our experiments, we use the AccessEval dataset [26], which comprises sets of queries spanning six real-world business domains: Education, Finance, Healthcare, Hospitality, Media, and Technology. Each set consists of:

### v2

@@ 2.1 Bias in LLMs
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of NLP tasks; however, they remain susceptible to representational bias and stereotyping. Prior research has shown that LLMs often absorb and reproduce harmful societal biases - especially those tied to spoken language [1], gender [32], race, religion, and health - across both contextual embeddings and generative outputs [6, 26, 13, 22, 2, 35]. These models frequently reflect the imbalances present in their training data, reproducing gendered, socioeconomic, and occupational assumptions rather than equitable or contextually informed representations [17, 28, 21, 33]. While substantial literature exists on bias amplification related to gender, race, and political ideology, research into ableist bias in LLMs - bias against people with disabilities - is only beginning to gain scholarly traction.
@@ 2.2 Disability bias
Emerging studies [12, 42] reveal that LLMs often encode and reflect deeply entrenched stereotypes about people with disabilities. For instance, models tend to associate disability-related terms with negative sentiment [18], echoing societal stigma found in training data. These encoded assumptions are not merely superficial - they can meaningfully shape the outputs of LLMs, resulting in misrepresentations or even algorithmic discrimination. This is particularly concerning in high-stakes contexts like hiring [19], where disclosure of a disability can lead to skewed responses. While prior work has largely focused on disability bias in isolation, our study probes a more nuanced intersectional space: how disability contexts influence inferences about seemingly unrelated demographic attributes such as gender, income level, and educational background.
@@ 2.4 Intersectionality of biases
There remains limited but growing attention on how intersecting identities influence LLM behavior. While several recent works [9, 18, 23, 24, 3, 34] have evaluated how multiple marginalized identities interact within prompts and input space, few studies have examined how disability cues lead to inferred biases about unrelated attributes. Notably, [23] explored ableist bias in tandem with other identity categories, using sentiment and social perception metrics. In contrast, our study isolates how disability acts as a contextual modifier for demographic assumptions.
To support such analysis, various benchmark datasets have emerged. For instance, [9] introduced a probing framework for identity-based behavior using persona-based prompts, while [24] presented a stereotype prevalence dataset for intersectional identities across LLMs. These newer datasets emphasize richer, sentence-level stimuli that better reflect real-world complexity [8, 11].
@@ 3.1 Dataset Selection
For our experiments, we use the AccessEval dataset [30, 31], which comprises sets of queries spanning six real-world business domains: Education, Finance, Healthcare, Hospitality, Media, and Technology. Each set consists of: